solve the following differential equations
1- (y-2)dx-(x-y-1)dy=0
2- (x-4y-9)dx+(4x+y-2)dy=0
1
Expert's answer
2019-11-25T11:46:19-0500

1) "(y-2)dx-(x-y-1)dy=0"

"\implies dy/dx=(y-2)/(x-y-1)"

Let "x=u+h;y=v+k"

"\implies dx=du ; dy=dv"

"dy\/dx=[(v+k)-2]\/[(u+h)-(v+k)-1)]"

"dv\/du=[v+(k-2)]\/[u-v+(h-k-1)]"


Now;

"k-2=0" ;

"h-k-1=0"

Solving these equations;

"k=2; h=3"


"\implies x=u+3; y=v+2"

"\implies dv/du= v/(u-v)"


Let; "v/u=z \implies v=uz"

"\implies dv/du=z+u(dz/du)"


"\implies z+u(dz/du)= z/(1-z)"

"\implies \int (z^{-2}-z^{-1})dz=\int du/u"

"\implies ln(uz)=-(1/z)+c"


Substituting the values of "u,v, z, h, k;" we get;


"(y-2)=ke^{-[(x-3)\/(y-2)]}" (Answer)



2) "(x-4y-9)dx+(4x+y-2)dy=0"

"\implies dy/dx=(4y-x+9)/(4x+y-2)"

Let "x=u+h;y=v+k"

"\implies dx=du ; dy=dv"

"dy\/dx=[4(v+k)-(u+h)+9]\/""[4(u+h)+(v+k)-2)]"

"dv\/du=[4v-u+(4k-h+9)]\/""[4u+v+(4h+k-2)]"


Now;

"4k-h+9=0" ;

"4h+k-2=0"

Solving these equations;

"k=-2; h=1"


"\implies x=u+1; y=v-2"

"\implies dv/du= (4v-u)/(4u+v)"


"v/u=z \implies v=uz"

"\implies dv/du=z+u(dz/du)"


"\implies z+u(dz/du)= (4z-1)/(z+4)"

"\implies \int [(z+4)/(z^{2}+1)]dz=-\int du/u"

"\implies 1/2(ln(u^2(z^2+1))=-tan^{-1}z+c"


Substituting the values of "u,v, z, h, k" ; we get;


"(x-1)^2+(y+2)^2=ke^{-8tan^{-1}[(y+2)\/(x-1)]}" (Answer)